OpenStudy (anonymous):

y=3x^2-6x+4 Can someone help me graph the parabola,find its vertex, x and y intercept

OpenStudy (anonymous):

Don't remember all the steps? =)

OpenStudy (anonymous):

The vertex i think i got it....(-1,7)

OpenStudy (anonymous):

almost

OpenStudy (anonymous):

x: -b/2a The b is already negative, so since b is (-6), then it makes it positive.

OpenStudy (anonymous):

-9

OpenStudy (anonymous):

So the vertex is -9,185 ??????

OpenStudy (anonymous):

mmm. ok, hold on.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

its a -3 not a postive 3

OpenStudy (anonymous):

oh ok, that changes some things.

OpenStudy (anonymous):

now what..how do i find the x and y axis

OpenStudy (triciaal):

when x = 0 what is y? when y = 0 what is x? y = ax^2 + bx + c identify a, b and c in this eqn 3x^2 -6x +4 vertex = -b/2a when a >= 0 it opens up; when a < 0 it turns down for this eqn is a >=0 or <0 use the points to sketch the graph